OpenStudy (anonymous):

Integrate 60/(3x-1)^.5

OpenStudy (anonymous):

is the exponent is 0.5?

OpenStudy (anonymous):

Yessir

OpenStudy (anonymous):

u can try to change 0.5 into 1/2 then make it as radical or square root

hartnn (hartnn):

use the substitution u= 3x-1 then du =... ?

OpenStudy (anonymous):

sir can u tell me the answer to this prob? i am not sure with my answer.

hartnn (hartnn):

sorry dear, this is not an answering site, it is learning/studying site. so if you want to learn how to do it, then we can proceed. but since you have tried it, can you show your steps ? i'll help you find the error if there is any.

OpenStudy (anonymous):

oh ok w8

OpenStudy (anonymous):

i use the formula of \[\int\limits_{?}^{?}\frac{ 1 }{ u } du = \ln u + c\]

hartnn (hartnn):

you won't need that formula

hartnn (hartnn):

what you got after plugging in u = 3x-1 ?

hartnn (hartnn):

and what you get as du = ... ?

OpenStudy (anonymous):

is it possible to use that formula sir? my du is \[\frac{ 1 }{ 2\sqrt{3x-1} }\]

hartnn (hartnn):

actually u was 3x-1 , not sqrt 3x-1 so, u =3x-1 gives du = 3dx makes sense ?

hartnn (hartnn):

and no, we don;t need log formula for this problem

OpenStudy (anonymous):

so we can simplify it into \[\int\limits_{?}^{?}\frac{ 1 }{ u ^{\frac{ 1 }{ 2 }} }\] ?

hartnn (hartnn):

yes, don't forget about the constants 60 and 3

hartnn (hartnn):

du = 3dx so dx = du/3 so we have 60/3 outside that integral

OpenStudy (anonymous):

yup, but im still confused on 1/u

hartnn (hartnn):

and for that integral you can directly use the formula for integral of x^n as 1/u^1/2 is just u^-1/2

OpenStudy (anonymous):

sir is the answer is \[40\sqrt{3x-1}+c\] ?

OpenStudy (anonymous):

i never thought that - exponent can be solved in the normal process. T_T

hartnn (hartnn):

yes, that answer is correct :) good!

OpenStudy (anonymous):

thank you very much sir. God bless po. :))

hartnn (hartnn):

you're welcome ^_^ God bless you too :)